OpenStudy (anonymous):

what is the value of x, when 10(x+2)=5(x+8)?

OpenStudy (solomonzelman):

\(\large\color{blue}{ 10(x+2)=5(x+8) }\) \(\large\color{blue}{ 5\times 2(x+2)=5(x+8) }\) \(\large\color{blue}{ 5(2x+4)=5(x+8) }\) \(\large\color{blue}{ 2x+4=x+8 }\) \(\large\color{blue}{ 2x-x=8-4 }\)

OpenStudy (solomonzelman):

and you can do the last thing